# LangBridge: Interpreting Image as a Combination of Language Embeddings

### v2

@@ Abstract
Recent years have witnessed remarkable advances in Large Vision-Language Models (LVLMs), which have achieved human-level performance across various complex vision-language tasks. Following LLaVA’s paradigm, mainstream LVLMs typically employ a shallow MLP for visual-language alignment through a two-stage training process: pretraining for cross-modal alignment followed by instruction tuning. While this approach has proven effective, the underlying mechanisms of how MLPs bridge the modality gap remain poorly understood. Although some research has explored how LLMs process transformed visual tokens, few studies have investigated the fundamental alignment mechanism. Furthermore, the MLP adapter requires retraining whenever switching LLM backbones. To address these limitations, we first investigate the working principles of MLP adapters and discover that they learn to project visual embeddings into subspaces spanned by corresponding text embeddings progressively. Based on this insight, we propose LangBridge, a novel adapter that explicitly maps visual tokens to linear combinations of LLM vocabulary embeddings. This innovative design enables pretraining-free adapter transfer across different LLMs while maintaining performance. Our experimental results demonstrate that a LangBridge adapter pre-trained on Qwen2-0.5B can be directly applied to larger models such as LLaMA3-8B or Qwen2.5-14B while maintaining competitive performance. Overall, LangBridge enables interpretable vision-language alignment by grounding visual representations in LLM vocab embedding, while its plug-and-play design ensures efficient reuse across multiple LLMs with nearly no performance degradation. See our project page at https://jiaqiliao77.github.io/LangBridge.github.io/.
Equal contribution.
@@ 1 Introduction
As visualized in Figure 2 through circular graphs where nodes represent different training phases spanning from early to late stages (e.g. 1- 4), we can draw two key conclusions: 1. Visual embeddings demonstrate strong correlations with semantically-related text tokens. As shown in Figure 2, green apple image patch exhibits high similarity with text tokens such as “Green Apple”. Similarly, the silhouetted sunset scene correlates with concepts like “Eclipse Sun”. This suggests that MLPs effectively map visual features into the text embedding space of LLMs, particularly in regions close to the semantic spaces formed by the corresponding text embeddings. 2. The projection capability of MLPs develops progressively during training. Figure 2 illustrates this evolution through multiple semantic pathways: from “numerous” to ”five”, from ”Arts” to “Green Apple”, and in the right scene, from “We Plat” to “Shakespeare”, from “Kiss” to “Wed Kiss”. These relationships become increasingly well-defined through the training stages, confirming our hypothesis that MLPs gradually learn to map visual features into regions proximate to their corresponding text embedding subspaces.
@@
Figure 2: Progressive semantic alignment in MLP adapters across training stages. Circular graphs demonstrate the evolution of visual-text token associations through four training phases (Pretrain-100, Pretrain-1000, Pretrain-2000, and Final-SFT). (a) For a bunch of apples image, the MLP progressively refines associations from the meaningless text (“numerous”, “Arts”) to meaningful text (“five”, “Green Apple”). (b) For a sunset silhouette scene and poem, semantic mappings evolve from the meaningless text (“We Plat”, “Kiss”) to contextually relevant tokens (“Shakespeare”, “Wed Kiss”), illustrating the MLP’s increasing capability to project visual features into LLM’s text embedding space.
@@ 2.1 Visual encoders in LVLMs
In recent years, large language models (LLMs) [50, 23, 25] have achieved remarkable advances in natural language processing. However, their inability to directly process visual information limits their effectiveness in multimodal tasks. To overcome this limitation, visual encoders [28, 47, 34, 27, 29] have been integrated, serving as the “eyes” of large vision-language models to leverage LLMs’ extensive knowledge [45, 26, 38] and strong instruction-following capabilities. Visual models that learn representations using only visual signals, such as DINOv2 [27] and diffusion models [29], often face challenges in achieving high performance, while models like CLIP [28] and SigLIP [47] align visual and text modalities effectively by training on extensive image-text datasets using contrastive learning. Recent research [41, 42, 36, 35] has also advocated for combining multiple visual encoders to better capture complex visual information. Nonetheless, aligning visual encoders with large language models remains challenging [44] due to differences in their representational spaces.
@@ 2.2 Connectors in LVLMs
To merge the visual extraction strengths of visual encoders with LLMs’ broad knowledge and instruction-following skills, researchers have explored methods to connect these components [54, 37, 43, 33, 48, 8, 9, 22, 4, 51, 6, 2, 10, 16, 52]. Pioneering work represented by models such as Flamingo [1], which uses cross-attention mechanisms, and BLIP-2 [13], which introduces the Q-former architecture, marks significant progress in this field. Subsequently, LLaVA [19, 18] adopted a two-stage learning approach, using a multilayer perceptron to map visual information into language embedding space, establishing itself as a most widely applicable alignment method [53, 21, 15, 5]. Departing from previous approaches, we propose an innovative mapping technique. Rather than mapping visual information directly to language embedding space, we introduce “language basis vectors” to learn linear combination weights of these basis vectors, creating a new way to represent visual information.
@@ 5.1 Experiment setup
Our architecture is composed of four key components: language model, visual encoder, connector, and language base vector. We selected LLaMA3-8B-Instruct [7], Qwen2-0.5B/7B-Instruct [39] and Qwen2.5-7B/14B-Instruct [40] as our language models. For brevity, we omit the “-Instruct” suffix in the remainder of this paper. We employed the most popular CLIP-ViT-L/14@336px[28] for the visual encoder. The connector is based on an MLP architecture with a hidden size of 5120. Regarding the language base vector, we carefully selected 19,200 high-frequency language tokens. Following the LLaVA [18] setup, we utilized LLaVA-CC3M-Pretrain-595K and LLaVA-v1.5-mix665k as the pretraining and instruction-tuning datasets, respectively, training for one epoch. Specific hyperparameter settings can be found in Appendix.
@@ 6 Conclusions
  • Discussion of current limitations and future directions S-C.
@@
  • POPE [14]: Evaluates the model’s ability to identify specific objects in images, aiming to detect object-level hallucinations. It uses ”yes/no” questions based on object annotations, with 50% of queries targeting existing objects and 50% targeting non-existing objects, employing random, popular, and adversarial sampling strategies.
@@
S-C Limitations

### v3

@@ Abstract
Recent years have witnessed remarkable advances in Large Vision-Language Models (LVLMs), which have achieved human-level performance across various complex vision-language tasks. Following LLaVA’s paradigm, mainstream LVLMs typically employ a shallow MLP for visual-language alignment through a two-stage training process: pretraining for cross-modal alignment followed by instruction tuning. While this approach has proven effective, the underlying mechanisms of how MLPs bridge the modality gap remain poorly understood. Although some research has explored how LLMs process transformed visual tokens, few studies have investigated the fundamental alignment mechanism. Furthermore, the MLP adapter requires retraining whenever switching LLM backbones. To address these limitations, we first investigate the working principles of MLP adapters and discover that they learn to project visual embeddings into subspaces spanned by corresponding text embeddings progressively. Based on this insight, we propose LangBridge, a novel adapter that explicitly maps visual tokens to linear combinations of LLM vocabulary embeddings. This innovative design enables pretraining-free adapter transfer across different LLMs while maintaining performance. Our experimental results demonstrate that a LangBridge adapter pre-trained on Qwen2-0.5B can be directly applied to larger models such as LLaMA3-8B or Qwen2.5-14B while maintaining competitive performance. Overall, LangBridge enables interpretable vision-language alignment by grounding visual representations in LLM vocab embedding, while its plug-and-play design ensures efficient reuse across multiple LLMs with nearly no performance degradation. See our project page at https://curryx-001.github.io/LangBridge.github.io/.
Equal contribution. 🖂
@@ 1 Introduction
As visualized in Figure 2 through circular graphs where nodes represent different training phases spanning from early to late stages (e.g. 1- 4), we can draw two key conclusions: 1. Visual embeddings demonstrate strong correlations with semantically-related text tokens. As shown in Figure 2, green apple image patch exhibits high similarity with text tokens such as “Green Apple". Similarly, the silhouetted sunset scene correlates with concepts like “Eclipse Sun". This suggests that MLPs effectively map visual features into the text embedding space of LLMs, particularly in regions close to the semantic spaces formed by the corresponding text embeddings. 2. The projection capability of MLPs develops progressively during training. Figure 2 illustrates this evolution through multiple semantic pathways: from “numerous" to "five", from "Arts" to “Green Apple", and in the right scene, from “We Plat" to “Shakespeare", from “Kiss" to “Wed Kiss". These relationships become increasingly well-defined through the training stages, confirming our hypothesis that MLPs gradually learn to map visual features into regions proximate to their corresponding text embedding subspaces.
@@
Figure 2: Progressive semantic alignment in MLP adapters across training stages. Circular graphs demonstrate the evolution of visual-text token associations through four training phases (Pretrain-100, Pretrain-1000, Pretrain-2000, and Final-SFT). (a) For a bunch of apples image, the MLP progressively refines associations from the meaningless text (“numerous", “Arts") to meaningful text (“five", “Green Apple"). (b) For a sunset silhouette scene and poem, semantic mappings evolve from the meaningless text (“We Plat", “Kiss") to contextually relevant tokens (“Shakespeare", “Wed Kiss"), illustrating the MLP’s increasing capability to project visual features into LLM’s text embedding space.
@@ 2.1 Visual encoders in LVLMs
In recent years, large language models (LLMs) [50, 23, 25] have achieved remarkable advances in natural language processing. However, their inability to directly process visual information limits their effectiveness in multimodal tasks. To overcome this limitation, visual encoders [28, 47, 34, 27, 29] have been integrated, serving as the “eyes" of large vision-language models to leverage LLMs’ extensive knowledge [45, 26, 38] and strong instruction-following capabilities. Visual models that learn representations using only visual signals, such as DINOv2 [27] and diffusion models [29], often face challenges in achieving high performance, while models like CLIP [28] and SigLIP [47] align visual and text modalities effectively by training on extensive image-text datasets using contrastive learning. Recent research [41, 42, 36, 35] has also advocated for combining multiple visual encoders to better capture complex visual information. Nonetheless, aligning visual encoders with large language models remains challenging [44] due to differences in their representational spaces.
@@ 2.2 Connectors in LVLMs
To merge the visual extraction strengths of visual encoders with LLMs’ broad knowledge and instruction-following skills, researchers have explored methods to connect these components [54, 37, 43, 33, 48, 8, 9, 22, 4, 51, 6, 2, 10, 16, 52]. Pioneering work represented by models such as Flamingo [1], which uses cross-attention mechanisms, and BLIP-2 [13], which introduces the Q-former architecture, marks significant progress in this field. Subsequently, LLaVA [19, 18] adopted a two-stage learning approach, using a multilayer perceptron to map visual information into language embedding space, establishing itself as a most widely applicable alignment method [53, 21, 15, 5]. Departing from previous approaches, we propose an innovative mapping technique. Rather than mapping visual information directly to language embedding space, we introduce “language basis vectors" to learn linear combination weights of these basis vectors, creating a new way to represent visual information.
@@ 5.1 Experiment setup
Our architecture is composed of four key components: language model, visual encoder, connector, and language base vector. We selected LLaMA3-8B-Instruct [7], Qwen2-0.5B/7B-Instruct [39] and Qwen2.5-7B/14B-Instruct [40] as our language models. For brevity, we omit the “-Instruct" suffix in the remainder of this paper. We employed the most popular CLIP-ViT-L/14@336px[28] for the visual encoder. The connector is based on an MLP architecture with a hidden size of 5120. Regarding the language base vector, we carefully selected 19,200 high-frequency language tokens. Following the LLaVA [18] setup, we utilized LLaVA-CC3M-Pretrain-595K and LLaVA-v1.5-mix665k as the pretraining and instruction-tuning datasets, respectively, training for one epoch. Specific hyperparameter settings can be found in Appendix.
@@ 6 Conclusions
  • Additional experimental results on a broader range of benchmarks for LLaVA-Next integrated with LangBridge. S-C
  • Discussion of Computational Cost. S-D
  • Discussion of current limitations and future directions S-E.
@@
  • POPE [14]: Evaluates the model’s ability to identify specific objects in images, aiming to detect object-level hallucinations. It uses "yes/no" questions based on object annotations, with 50% of queries targeting existing objects and 50% targeting non-existing objects, employing random, popular, and adversarial sampling strategies.
@@
S-C More experiments results
To further evaluate the generalizability and robustness of our method, we integrate it into LLaVA-Next. Our model first uses a Qwen2-0.5B pretrained LangBridge module, which is then combined with Qwen2-7B for further supervised fine-tuning (SFT). We compare it to a standard MLP baseline with same backbone and evaluate across a range of benchmarks. As shown in Table S-2, our model achieves consistent improvements over the baseline in fine-grained image analysis tasks, including AI2D, ChartQA, and DocVQA, while maintaining parity on MMVP. For grounding and real-world reasoning tasks (Table S-3), our approach significantly outperforms the baseline on RefCOCO (+17.06%) and RefCOCO+ (+15.23%), and performs comparably on RefCOCOg and MMRealWorld. In the video understanding domain (Table S-4), the model improves results on Seed-Video and Seed2-Video while maintaining similar performance on VideoMME and MMT. Lastly, as shown in Table S-5, our method improves accuracy on TextVQA, GQA, and MME, demonstrating its effectiveness on widely used benchmarks, with only a slight drop on MMMU. These results collectively demonstrate that LangBridge consistently matches or outperforms standard MLPs across a wide range of tasks, showcasing strong generalization capability.
Table S-2: Results on Fine-grained Benchmarks
Method AI2D ChartQA DocVQA MMVP
Baseline 0.763 0.8632 0.741 43.3
Our Model 0.766 (100.39%) 0.877 (101.60%) 0.758 (102.29%) 43.3 (100.00%)
Table S-3: Results on Grounding and Real-World Benchmarks
Method RefCOCO RefCOCO+ RefCOCOg MMRealWorld
Baseline 0.387 0.348 0.6377 0.4405
Our Model 0.453 (117.06%) 0.401 (115.23%) 0.622 (97.53%) 0.4436 (100.70%)
Table S-4: Results on Video Benchmarks
Method VideoMME Seed-Video Seed2-Video MMT
Baseline 50.77 0.444 0.494 54.43
Our Model 50.71 (99.88%) 0.450 (101.35%) 0.504 (102.03%) 53.98 (99.17%)
Table S-5: Results on Commonly Used Datasets
Method TextVQA GQA MME MMMU
Baseline 0.654 0.651 1886 0.4356
Our Model 0.665 (101.68%) 0.652 (100.15%) 1928 (102.23%) 0.4178 (95.91%)
S-D Computational Cost
We follow LLaVA-Next settings and conduct experiments on H100 using its SFT data. As shown in the table below, LangBridge incurs only a  10% increase in training time compared to the baseline, while significantly reducing pre-training time.
Vision Backbone LLM MLP Speed (s/iter) Time (h) Parameters Data Hardware
[…]
S-E Limitations